OpenStudy (anonymous):

Solve the equation for a: a a/2+ a/3= 5 . A)2 B)6 C)3 D)5 E)30

OpenStudy (anonymous):

y wont this thing do fractions

OpenStudy (texaschic101):

it will.... 2/3....do it like that

OpenStudy (wmj259):

\[\frac{ 2 }{ 3 } + a =5?\]

OpenStudy (wmj259):

oh.

OpenStudy (wmj259):

try to find a least common denominator for all the terms. And then you can isolate the a.

OpenStudy (texaschic101):

a/2 + a/3 = 5 --- multiply everything by the common denominator of 6 3a + 2a = 30 5a = 30 a = 6

OpenStudy (wmj259):

LCD would be 30.

OpenStudy (anonymous):

what?? is that even english?? lol i dont speak math can you dumb it down for me please

OpenStudy (anonymous):

my computer just froze

OpenStudy (texaschic101):

if you have an equation with fractions, and you want to get rid of the fractions, multiply everything in the equation by the common denominator. a/2 + a/3 = 5 (or 5/1).....now the common denominator of 2,3, and 1 is 6. So multiply everything by 6. 6(a/2) + 6(a/3) = 6(5) -- just to let you know, a/2 is the same as (1/2)a and a/3 is the same as (1/3)a 6(1/2)a + 6(1/3)a = 6(5) 6/2a + 6/3a = 30 3a + 2a = 30 5a = 30 -- divide both sides by 5 5a/5 = 30/5 a = 6